If the measure of the third angle of the triangle is 35 degrees more than three times the measure of either of the other two angles, find the measure of each triangle.
Answer by rfer(12657)

x+x+3x+35=180
5x=180-35
5x=145
x=145/5
x=29 for two
3x+35=122